OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. Integral symbol 1(at top) -1 (at bottom) (3x^2 +x)e^-x dx So I know this is the tabular method, and so far I have : u=(3x^2+x) derivative of U: Anti-deriv. of Dv: 3x^2+x e^-x =6x -e^-x =6 e^-x =0 I dont know how to do the rest!

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\frac{3x^2+x}{e^x} dx\] Select 3x^2+x as your u and e^-x as your dv u = 3x^2+x du = 6x + 1 dx (take the derivative to find du) v = -e^-x (take the antiderivative to find v) dv = e^-x IBP states that... \[\int\limits u dv = uv - \int\limits v du\] Using our values for u, du, v and dv, we find that... \[\frac{-(3x^2+x)}{e^x} -\int\limits \frac{-(6x+1)}{e^x} dx\] You will see we will have to use IBP once more, but we are getting closer to the solution. u = 6x+1 du = 6 dx v = e^-x dv = -e^-x dx du = 6x + 1 dx (take the derivative to find du) v = -e^-x (take the antiderivative to find v) dv = e^-x Now... \[\frac{-(3x^2+x)}{e^x}-(\frac{6x+1}{e^x}- 6\int\ \frac{1}{e^x}dx)\] You will find through the integration of the final e^-x and a bit of algebra that... \[\int\limits\frac{3x^2+x}{e^x} dx=\frac{-(3x^2+7x+7)}{e^x} + C \] Now using this, we can find the definite integral Evalutate -(3x^2+7x+7)e^-x at 1, then take that value and subtract away -(3x^2+7x+7)e^-x evaluated at -1 That should give you -17/e + 3e which is about 1.900894985...

OpenStudy (anonymous):

Thank you so much!!! Just one question, how do you know how many times you have to use IBP?

OpenStudy (anonymous):

Until the integrand has been simplified to a point where you can integrate without IBP. I stopped using IBP when I got to the integration of 1/e^x or e^-x because its a fairly simple integration problem, being -e^-x